OpenStudy (anonymous):

The ordered pairs (1, 3), (2, 9), (3, 27), (4, 81), and (5, 243) represent a function. What is a rule that represents this function? And The ordered pairs (1, 25), (2, 36), (3, 49), (4, 64), and (5, 81) represent a function. What is a rule that represents this function? Thanks guys!

OpenStudy (anonymous):

(x,y) x=n y=3(n_last)

OpenStudy (anonymous):

Shouldnt it be more like y = x3 or y = 3x for the first one?

OpenStudy (amistre64):

yes, the first one represents powers of 3\[(x,3^x)\] the second one looks a little square to me.....

OpenStudy (anonymous):

Is the second one y=x^5 maybe?

OpenStudy (amistre64):

no, give me a list of perfect square from 1 to 9 ...

OpenStudy (amistre64):

1^2 = 1 2^2 = 4 3^2 = 9 ...

OpenStudy (anonymous):

Ok? You want me to continue that or no?

OpenStudy (amistre64):

continue it for the integers from 1 to 9; so from 1^2 to 9^2

OpenStudy (anonymous):

Oh ok, you already did the first three, so the next would be: 4^2=16 5^2=25 6^2=36 7^2=49 8^2=64 and 9^2=81

OpenStudy (amistre64):

now, notice that our outputs given are 5^2=25 6^2=36 7^2=49 8^2=64 9^2=81 now, when x=1, we need x to be 5 when x=2, we need x to be 6 when x=3, we need x to be 7 we need the x parts to be shifted by +4 do you see it?

OpenStudy (anonymous):

I think I get ya, but whats an equasion for that? I understand the problem now though=) thanks

OpenStudy (amistre64):

well it would be nice if x^2 would get us where we need to me, but we discovered the x needs to be shifted by +4 in order to make a math sooo y = (x+4)^2

OpenStudy (amistre64):

soo many typos, sooo little time lol

OpenStudy (anonymous):

Haha ok, I get it now...makes alot more sense=) and yes..thanks for your melp eben wif tybos=)

OpenStudy (amistre64):

just to clarify ;) it would be nice if x^2 would get us where we need to be, but we discovered that x needs to be shifted by +4 in order to make a match ... good luck

OpenStudy (anonymous):

Thanks again amistre, I really do appreciate it=D